OpenStudy (anonymous):

a particle moves along a line so that, at time t, its position is s(t)=8sin2t. for what values of t does the particle change direction?

OpenStudy (anonymous):

at the points of inflection?

OpenStudy (anonymous):

no wait, it's at the relative max and minimums right?

OpenStudy (anonymous):

yes

OpenStudy (agreene):

lol I was going to answer as if this was a displacement vector >.>

OpenStudy (anonymous):

its setting the first derivative = 0, and solving for t

OpenStudy (anonymous):

s'(t) = 8(sin2t)(cos2t)(2) = 16(sin2t)(cos2t)

OpenStudy (anonymous):

sin2t = 0 when t = pi/2 cos2t = 0 when t = pi

OpenStudy (anonymous):

i did a similar problem, i forgot roflmao

OpenStudy (anonymous):

how would i solve for t?

OpenStudy (anonymous):

sin is zero at pi + npi where n is an integer

OpenStudy (anonymous):

the answer is tan = pi/4

OpenStudy (anonymous):

is there a restriction for t?

OpenStudy (anonymous):

um the answer states: t=pi/4 + pik,3pi/4 for positive integers k

OpenStudy (anonymous):

-4cos2t+sin2t=0 -4(1-2sin^2t)(2sintcost)=0 -4(2sintcost-4sin^3tcost)=0 -8sintcost+16sin^3tcost=0 2sin^3tcost=sintcost sin^2t=1/2 sint=sqrt2/2 t=45,135

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

nope hahah

OpenStudy (anonymous):

that was a question u answered, 13 days ago.

OpenStudy (anonymous):

yeah i'm trying to find that hw problem, but that guy didn't help

OpenStudy (anonymous):

he was working in degrees or something

OpenStudy (anonymous):

and using identities i think

OpenStudy (anonymous):

degrees u can easily convert to radians.

OpenStudy (anonymous):

Shouldn't it be 2t=pi/2

OpenStudy (anonymous):

yea, but the answer is in tan, and the answer he gave me is both in cos and sin?

OpenStudy (anonymous):

yeah, my answer is incomplete

OpenStudy (anonymous):

first can u explain how u came up with the derivative.

OpenStudy (anonymous):

b/c the derivative i got was s'(t)=16sin2t

OpenStudy (anonymous):

16cos2t*

OpenStudy (anonymous):

that's right yeah

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

i got 16cos2t, not 16sin2tcos2t

OpenStudy (anonymous):

waitwait

OpenStudy (anonymous):

i almost have it, hang on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you sure it changes direction at the relative max/min? I wanna say it changes direction at points of inflection..

OpenStudy (anonymous):

nvm, I thought I had it, gotta brush up on my trig

OpenStudy (anonymous):

um ok